# Agentic Large Language Models, a survey

### v2

@@ 3.3 Assistants
The progress in reasoning and decision making has improved the accuracy and usability of LLMs for everyday tasks. Also, LLMs can act through their use of tools. Tool-enabled LLMs can be used as virtual assistants. There is commercial interest in this element of agentic AI, and much activity has been reported. Furthermore, assisting humans generates more training data, on which LLMs can be pretrained and finetuned.
@@
We start our review with conversational assistants, and then continue to shopping, travel, medical, and financial trading support. An assistant can be seen here as a use-case of an agentic LLM for a specific range of tasks or a specific working domain.
3.3.1 Conversational Assistants
@@
Agentic LLMs can be used to make Human-Computer interaction more natural (Neszlényi et al., 2024; Oluwagbade, 2024). The AssistantGPT system supports a diverse range of operations, including web searches, API interactions via OpenAPI schemas, voice conversations, and command execution through the shell. The system consists of an LLM with access to tools, a planner, and memory (see Figure 16). The system is designed for deployment in an educational setting, a corporate setting, and to support remote work environments such as Teams and Slack.
@@
Cabrero-Daniel et al. (2024) describe how LLM meeting assistants can improve agile software development team meetings, to generate favorable results for preparation and live assistance during Scrum meetings, although some testers remarked that LLM interventions felt unnatural and inflexible.
@@
A system to facilitate group conversations is MUCA (Mao et al., 2024), supporting What, When and Who questions, consisting of a sub-topic generator, dialog analyzer, and conversational strategies arbitrator. Wei et al. (2024b) report improved collaboration through the use of LLM agents in a collaborative learning classroom setting. Another study reports improved work efficiency in a collaborative task scheduling experiment (Bastola et al., 2023).
@@
A different type of assistant is the thinking assistant. This assistant tries to improve (human) reflective thinking for difficult decisions, by asking instead of answering (Park and Kulkarni, 2023).
@@
Conversational assistants have mostly grown out of regular LLMs, sometimes finetuned for a particular area of expertise or domain. Some approaches use a specialized multi-LLMs approach, specializing LLMs for different sub-tasks.
@@
Basic LLMs generally lack inherent knowledge of e-commerce concepts. Jin et al. (2024b) created the Multi task Online Shopping Benchmark. Shopping MMLU consists of 57 tasks covering 4 major shopping skills: concept understanding, knowledge reasoning, user behavior alignment, and multi-linguality. Vedula et al. (2024) provide question suggestion for shopping assistants based on product metadata. ChatShop presents a information-seeking centric evaluation of language agents (Chen et al., 2024a). Zhang et al. (2024a) created an E-commerce shopping assistant named LLaSa. They created an instruction dataset comprising 65,000 samples and diverse tasks, and trained the model through instruction tuning. The system scores high on the ShopBench benchmark.
@@
Related to shopping assistants are assistants for booking flights. A flight booking assistant based on LLaMa 2 and RAG has been developed (Manasa et al., 2024). In user testing the system scored positive in understanding user preferences and efficient completion of the booking process.
@@
In specialized domains, operations support assistant have been developed. For example, to automate flight planning under wind hazards (Tabrizian et al., 2024), or for flight arrival scheduling (Zhou et al., 2024b). Wassim et al. (2024) introduce Drone-as-a-Service operations from text user requests. As agentic LLM technology matures, more specialized domain assistants will be developed.
3.3.2 Medical Assistants
@@
The interest of the field of medicine in LLMs is large (Thirunavukarasu et al., 2023; Clusmann et al., 2023; Mehandru et al., 2024). A recent study showed LLMs scoring higher on diagnoses than trained doctors (Goh et al., 2024). In medical conversation for medical note generation, LLMs are also exceeding the performance of human scribes (Yuan et al., 2024a). Another study finds similar results, but also points to shortcomings in specific areas (Panagoulias et al., 2024).
@@
Sudarshan et al. (2024) report on an experiment with an agentic workflow for generating patient-friendly medical reports, using the Reflexion approach (Section 2.2.2, (Shinn et al., 2024)), to comply with the 21th Century Cures Act that grants patients the right to access their health record data.
@@
A study by Qiu et al. (2024b) reports a wealth of opportunities for LLMs in medicine, ranging from clinical workflow automation to multi-agent aided diagnosis. Ullah et al. (2024) provide a scoping review on the use of ChatGPT for diagnostic medicine. Their main conclusion is that medical and ethical knowledge is necessary when training and finetuning these models. A challenge for the adoption of LLM in medicine are concerns about the quality, accuracy, and the comprehensiveness of LLM-generated answers. Das et al. (2024) describe how to mitigate common pitfalls such as hallucinations, incoherence, and lost-in-the-middle problems. They do so by implementing an agentic architecture, changing the LLM’s role from directly generating answers, to that of a planner in a retrieval system. The LLM-agent orchestrates a suite of specialized tools that retrieve information from various sources.
@@
In the domain of medical education, Wei et al. (2024a) use a multi agent framework to create copilots that emulate extensive real-world medical training environments (see Figure 17).
@@
A benchmark for retrieval-augmented generation in the medical domain is (Qiao et al., 2024).
3.3.3 Trading Assistants
A further specialized domain is financial trading. Already many algorithms are used in financial organizations for trading decisions. The interest in agentic LLMs in the financial world is large (Ding et al., 2024).
InvestorBench is a benchmark for financial trading systems (Li et al., 2024a). FinAgent is a tool-augmented multimodal agent for financial trading (Zhang et al., 2024b). It contains a market intelligence module, which is able to extract insights from multi-modal datasets of asset prices, visual representations, news, and expert analyses. The system can also perform query retrieval, and performs reflection in a low-level module for technical analysis, and in a high-level module to analyze past trading decisions.
FinRobot is an agentic LLM for financial analysis, to assist human traders. (Yang et al., 2024a). FinRobot can provide document analysis and generation, and market forecasts for individual stocks. FinMem is an agentic LLM framework devised for financial decision-making (Yu et al., 2024b). It features a layered memory system and adjustable character design for the trading agent. FinMem is inspired by the generative agents framework by Park et al. (2023) (see Section 4.3).
So far, most financial market machine learning has focused on single agent systems. An approach called TradingAgents uses a multi agent system to replicate real-world trading firms’ collaborative dynamics (Xiao et al., 2024). TradingAgents simulates LLM-powered agents in specialized roles such as fundamental analysts, sentiment analysts, technical analysts, and traders with varied risk profiles. The outcome of the system is a buy or sell advice to a human manager. A simulation showed that it outperformed baseline models.
3.3.4 Science Assistants
@@
The workflow of scientific experimentation is relatively standardized in certain fields of science. For example, in machine learning, ideas are generated, a hypothesis is formulated, an experiment is designed, datasets are acquired, experiments are performed, results are interpreted and a report is produced. This workflow has attracted researchers to experiment with agentic LLMs, see Eger et al. (2025) for a survey.
@@
AI Scientist (Lu et al., 2024a) is a framework to automate the process of scientific discovery, from idea generation to paper writing, including a review process. Users must specify a topic, and provide an experimentation template and indicate datasets. The authors report experiments in three areas of machine learning: diffusion modeling, transformer-based language modeling, and learning dynamics, with promising results. To improve idea generation and reviewing, the agent accesses open sources. Current limitations include limited experiments, incorrect implementation of ideas, and visual errors when the paper is produced. Rarely are entire results hallucinated. Results of the AI scientist are recommended to be taken as hints of promising ideas, worthy of a follow up study (Lu et al., 2024a).
@@
SWE-Agent (short for software engineering agent) (Yang et al., 2024b) aims to automate the process of software discovery, to help agents to autonomously use computers to solve software engineering tasks. SWE-agent introduces tools to create and edit code files, navigate through software repositories, and execute computer programs. Experiments on coding benchmarks such as HumanEvalFix achieve high success rates of over 80%. This success is attributed to the interactive design of the agent (see Figure 18).
@@
MLGym (Nathani et al., 2025) follows the popular Gym reinforcement learning framework (Brockman et al., 2016). Gym provides a standardized interface between environment and agent. Introduced in 2016, it accelerated the development of reinforcement learning algorithms, facilitating progress in the field. Taking further inspiration from SWE-agent (such as file editing capabilities) the MLGym work shows how the process of scientific discovery can be modeled as an interactive process. Applications are discussed in fields ranging from data science, game theory, computer vision, reinforcement learning, to natural language processing. Experiments are reported with commercial LLMs (OpenAI, Meta, Google, and Anthropic).
Google and OpenAI have both released Deep Research agents. These agents can perform multi-step research tasks, synthesizing online information. They are built with a reasoning LLM and use retrieval augmentation for finding information sources. The systems are able to create scientific research papers that look impressive, but are often not error-free.
3.4 Discussion
@@
3.4.1 Grounding Actions in the Real World
For agents to act in the real world, their understanding must be grounded in the world. They should sense their surroundings, understand it, and take actions that make sense. LLMs that were only trained on a language corpus may suggest actions such as trying to open doors that do not exist, or moving kitchen items that are not present. World Models and VLAs provide a step towards this world understanding, so that robots and assistants can make actions that make sense.
World Models and Action Models
How can LLMs learn to act? Many model-based reinforcement learning approaches learn world models that are surrogates of the real environment, in which state transitions and state rewards are learned, so that the agent can learn a policy of optimal actions. These world models can be used to train LLMs. An alternative is to train Action models directly from robotic actions, in the form of combined Vision-Language-Action models. The goal is that a VLA allows robots to perform a natural language instruction by looking at a visual scene.
Grounded Action, Tools
@@ 3.4.1 Grounding Actions in the Real World
For LLMs to be able to take actions, they must be able to call tools. LLMs have been trained to interact with a wide variety of tool-APIs.
3.4.2 Use Case: Assistants
@@
Agentic LLMs is an active field of research, some of which is aimed at making assistants ready for commercial deployment. If they work well, there may be a large market for robotic assistants that perform tedious or dangerous work, and for LLM agents that outperform humans in, for example, medical and trading decisions. However, such commercial deployment is still some time into the future, also because important ethical and legal questions should be resolved. If an LLM assistant provides medical advice, and a patient suffers, who is responsible? If an assistant suggests a certain trade, and a trader loses a sum of money, who is liable? Also, the impact on society and the work force has economic implications. Further research is necessary to resolve these questions (Akata et al., 2020).
AI Scientist (Lu et al., 2024a), SWE-Agent (Yang et al., 2024b), and MLGym (Nathani et al., 2025) demonstrate that automated scientific discovery is increasingly feasible using agentic LLMs. These systems achieve promising results across tasks in machine learning, software engineering, and beyond, due in part to interactive, tool-augmented designs. Nonetheless, limitations persist in the form of incorrect implementations, visual errors, and occasional hallucinations. These agents should be used to generate promising leads rather than definitive results. Recent releases of “Deep Research” agents by Google and OpenAI further illustrate the rapid progress of automated science, although they too are not error-free.
4 Interacting
We will now turn to the third category of the survey: interacting agents. Traditional LLMs passively respond to user queries, have no memories for interaction histories beyond their context window, and do not plan future steps of interaction ahead. This is shifting with agentic LLMs: LLMs can have memories and planning abilities. Reflective loops can lead to actions at their own initiative. This opens new potential for social interaction with users and other machine agents.
We first briefly discuss social and interactive capabilities in traditional, non-agentic LLMs, to identify the roots of their ability to interact with users and agents. Second, we discuss pairs or small teams of agentic LLMs that have role-based interactions to complete a task, game, or experiment. Third, we turn to open-ended interactions of LLM agents, interacting semi-spontaneously without prior role assignment, forming LLM societies that show self-organizing behavior, social dynamics, and emergent norms.
4.1 Social Capabilities of LLMs

### v3

@@ 3.3. Assistants
The progress in reasoning and decision making has improved the accuracy and usability of LLMs for everyday tasks. Also, LLMs can act through their use of tools. Tool-enabled LLMs can be used as virtual assistants. The use of agentic LLMs as assistants has received commercial interest, and much activity has been reported. An additional advantage is that by assisting humans, the agents generate new training data on which the LLMs can be pretrained and finetuned.
@@
We start our review of assistants with conversational assistants, and then continue to shopping, travel, medical, and financial trading assistants. An assistant can be seen here as a use-case of an agentic LLM for a specific range of tasks or a specific working domain.
3.3.1. Conversational Assistants and Negotiation
@@
Agentic LLMs can be used to make Human-Computer interaction more natural (neszlenyi2024assistantgpt; oluwagbade2024conversational). The AssistantGPT system supports a diverse range of operations, including web searches, API interactions via OpenAPI schemas, voice conversations, and command execution through the shell. The system consists of an LLM with access to tools, a planner, and memory (see Figure 16). The system is designed for deployment in an educational setting, a corporate setting, and to support remote work environments such as Teams and Slack.
@@
cabrero2024exploring describe how LLM meeting assistants can improve agile software development team meetings, to generate favorable results for preparation and live assistance during Scrum meetings, although some testers remarked that LLM interventions felt unnatural and inflexible.
@@
A system to facilitate group conversations is MUCA (mao2024multi), supporting What, When and Who questions, consisting of a sub-topic generator, dialog analyzer, and conversational strategies arbitrator. wei2024improving report improved collaboration through the use of LLM agents in a collaborative learning classroom setting. Another study reports improved work efficiency in a collaborative task scheduling experiment (bastola2023llm).
@@
A different type of assistant is the thinking assistant. This assistant tries to improve (human) reflective thinking for difficult decisions, by asking instead of answering (park2023thinking).
Another kind of assistants are research agents, that gather information by combining information from different tools and knowledge sources based on context, and synthesize and summarize the feedback based on the information need. For instance, Vogetal2025 developed a REWOO-based process mining assistant that combines a process discovery agent to discover processes and inefficiencies with parallel process research agents, an optional human-in-the-loop to interpret results from a domain-specific real-world knowledge perspective, and a reporting agent to summarize results. Research agents are popular initial use cases. They are relatively low risk, as actions are limited to various tools to gather data and content, and they build on the strengths of LLMs to analyze intent and synthesize information rather than being the knowledge source itself.
@@
Conversational assistants have mostly grown out of regular LLMs, sometimes finetuned, grounded and customized for a particular area of expertise or domain. Some approaches use a specialized multi-LLMs approach, specializing LLMs for different sub-tasks.
@@
Basic LLMs generally lack inherent knowledge of e-commerce concepts. jin2024shopping created the Multi task Online Shopping Benchmark. Shopping MMLU consists of 57 tasks covering 4 major shopping skills: concept understanding, knowledge reasoning, user behavior alignment, and multi-linguality. vedula2024question provide question suggestion for shopping assistants based on product metadata. ChatShop presents evaluation focused on information-seeking (chen2024chatshop). zhang2024llasa created an E-commerce shopping assistant named LLaSa. They created an instruction dataset comprising 65,000 samples and diverse tasks, and trained the model through instruction tuning. The system scores high on the ShopBench benchmark.
Automated negotiation between agents has been studied extensively in AI (jonker2017automated). LLM based negotiation introduces a risk of unexpected bias. A study by kirshner2025talking notes a tendency towards reaching agreement, which may influence contract terms. A further experimental analysis finds risks of overspending or unreasonable deals for automated negotiation (zhu2025automated). Section 4.1.2 also discusses negotiation, in a social setting.
@@
Assistants for booking flights are related to shopping assistants. (manasa2024towards) have developed a flight booking assistant based on LLaMa 2 and RAG. In user testing the system scored positive in understanding user preferences and efficient completion of the booking process.
@@
In specialized domains, operations support assistant have been developed. For example, to automate flight planning under wind hazards (tabrizian2024using), and for flight arrival scheduling (zhou2024flight). wassim2024llm introduce Drone-as-a-Service operations from text user requests. As agentic LLM technology matures, more specialized domain assistants will be developed.
3.3.2. Medical Assistants
@@
The field of medicine has shown great interest in LLMs (thirunavukarasu2023large; clusmann2023future; mehandru2024evaluating). A recent study showed LLMs scoring higher on diagnoses than trained human doctors (goh2024large). In medical conversations, for medical note generation, LLMs are also exceeding the performance of human scribes (yuan2024continued). Another study finds similar results, but also points to shortcomings in specific areas (panagoulias2024evaluating).
@@
sudarshan2024agentic report on an experiment with an agentic workflow for generating patient-friendly medical reports, using the Reflexion approach (Section 2.2.2, (shinn2024reflexion)), to comply with the 21th Century Cures Act that grants patients the right to access their health record data.
@@
A study by qiu2024llm reports a wealth of opportunities for LLMs in medicine, ranging from clinical workflow automation to multi-agent aided diagnosis. ullah2024challenges provide a scoping review on the use of ChatGPT for diagnostic medicine. Their main conclusion is that medical and ethical knowledge is necessary when training and finetuning these models. A challenge for the adoption of LLM in medicine are concerns about the quality, accuracy, and the comprehensiveness of LLM-generated answers. das2024improved describe how to mitigate common pitfalls such as hallucinations, incoherence, and lost-in-the-middle problems. They do so by implementing an agentic architecture, changing the LLM’s role from directly generating answers, to that of a planner in a retrieval system. The LLM-agent orchestrates a suite of specialized tools that retrieve information from various sources.
@@
In the domain of medical education, wei2024medco use a multi agent framework to create copilots that emulate extensive real-world medical training environments (see Figure 17).
@@
A benchmark for retrieval-augmented generation in the medical domain is (qiao2024benchmarking).
3.3.3. Science Assistants
@@
The workflow of scientific experimentation is relatively standardized in certain fields of science. For example, in machine learning, ideas are generated, a hypothesis is formulated, an experiment is designed, datasets are acquired, experiments are performed, results are interpreted and a report is produced. Google and OpenAI have both released Deep Research agents. These agents can perform multi-step research tasks, synthesizing online information. They are built with a reasoning LLM and use retrieval augmentation for finding information sources. The systems are able to create papers that look impressive, but may contain errors, as also indicated by the accompanying disclaimers. This workflow has attracted researchers to experiment with agentic LLMs, see eger2025transforming for a survey.
@@
AI Scientist (lu2024ai) is a framework to automate the process of scientific discovery, from idea generation to paper writing, including a review process. Users must specify a topic, and provide an experimentation template and indicate datasets. The authors report experiments in three areas of machine learning: diffusion modeling, transformer-based language modeling, and learning dynamics, with promising results. To improve idea generation and reviewing, the agent accesses open sources. Current limitations include limited experiments, incorrect implementation of ideas, and visual errors when the paper is produced. Rarely are entire results hallucinated. Results of the AI scientist are recommended to be taken as hints of promising ideas, worthy of a follow up study (lu2024ai).
A similar approach was taken in Agent Laboratory, where the agentic system was positioned as a crew of research assistants, working guided by human researchers across literature review, experimentation, and report writing, and producing experiments, code repositories and a final report (schmidgall2025agentlaboratoryusingllm). The results were evaluated through a survey. Early human involvement was found to improve the quality of research. The authors claim that the generated code outperformed prior results, with a substantial reduction in research effort.
@@
SWE-Agent (short for software engineering agent) (yang2024swe) aims to automate the process of software discovery, to help agents to autonomously use computers to solve software engineering tasks. SWE-agent introduces tools to create and edit code files, navigate through software repositories, and execute computer programs. Experiments on coding benchmarks such as HumanEvalFix achieve high success rates of over 80%. This success is attributed to the interactive design of the agent (see Figure 18).
@@
MLGym (nathani2025mlgym) follows the popular Gym reinforcement learning framework (brockman2016openai). Gym provides a standardized interface between environment and agent. Introduced in 2016, it accelerated the development of reinforcement learning algorithms, facilitating progress in the field. Taking further inspiration from SWE-agent (such as file editing capabilities) the MLGym work shows how the process of scientific discovery can be modeled as an interactive process. Applications are discussed in fields ranging from data science, game theory, computer vision, reinforcement learning, to natural language processing. Experiments are reported with commercial LLMs (OpenAI, Meta, Google, and Anthropic).
Science assistants are moving from isolated simulations to human/agent research collaborations. gottweis2025aicoscientist describe an attempt to discover significant scientific knowledge and validate these findings in real world experiments in their AI Co-scientist work. Given a general research goal or idea, AI Co-scientist uses multiple agents for idea generation, reflection, ranking, evolution, proximity and meta-review, ultimately to generate research hypotheses and plans. Such collaborations were validated in real world laboratory experiments in drug repurposing, novel treatment targets and explaining specific mechanism in gene transfer evolution related to microbiological resistance (gottweis2025aicoscientist; He2025; Penades2025).
Another example of real world validation is The Virtual Lab. This approach employs a PI agent directing a crew of specialized agents in chemistry, computer science and bioinformatics tools, that collaborate with a human researcher to identify new SARS-CoV-2 nanobodies. Promising results in experimental validation were reported (Swanson2024).
3.3.4. Trading Assistants
Another important specialized domain is financial trading. Already many algorithms are used in financial organizations to support trading decisions. The interest in agentic LLMs in the financial world is large (ding2024large).
InvestorBench is a benchmark for financial trading systems (li2024investorbench). FinAgent is a tool-augmented multimodal agent for financial trading (zhang2024finagent). It contains a market intelligence module, which is able to extract insights from multi-modal datasets of asset prices, visual representations, news, and expert analyses. The system can also perform query retrieval, and performs reflection in a low-level module for technical analysis, and in a high-level module to analyze past trading decisions.
FinRobot is an agentic LLM for financial analysis, to assist human traders. (yang2024finrobot). FinRobot can provide document analysis and generation, and market forecasts for individual stocks. FinMem is an agentic LLM framework devised for financial decision-making (yu2024finmem). It features a layered memory system and adjustable character design for the trading agent. FinMem is inspired by the generative agents framework by park2023generative (see Section 4.3).
So far, most financial market machine learning has focused on single agent systems. An approach called TradingAgents uses a multi agent system to replicate real-world trading firms’ collaborative dynamics (xiao2024tradingagents). TradingAgents simulates LLM-powered agents in specialized roles such as fundamental analysts, sentiment analysts, technical analysts, and traders with varied risk profiles. The outcome of the system is a buy or sell advice to a human manager. A simulation showed that it outperformed baseline models.
3.4. Discussion
@@
We have surveyed the individual methods for world models and agentic assistants. In order to dig deeper into the capabilities of LLM assistants, we will now discuss systems to perform scientific research in more detail.
3.4.1. In Depth: AI Scientist
We have seen how AI Scientist (lu2024ai), AI Co Scientist (schmidgall2025agentlaboratoryusingllm) and SWE-Agent (yang2024swe) are used to perform a full scientific and software engineering workflow. Google and OpenAI have released Deep Research agents that produce research reports in half an hour’s time. Impressive results are being reported—ideas have been generated, datasets have been downloaded, experiments have been performed, and scientific articles have been written. Although the field is still young, we will analyze early results and discuss possible consequences for scientific research.
First, we note that tools such as Alphaxiv are already surprisingly useful for summarizing the content of scientific papers, and offer social community building tools, and LMnotebook can produce readable blogs and reports that summarize the essence of scientific papers remarkably well. Since (1) LLMs are good at text generation, (2) agents can access tools to perform external tasks such as running machine learning experiments, and, thanks to self reflection, (3) agents can learn from their mistakes and try again, it should come as no surprise that certain scientific workflows are a good fit for agentic LLMs. The elements for performing basic scientific experiments and reporting about them are in place.
To what extent are LLM science tools able to perform independent, creative, high quality, research? Tools such as AI Scientist require prompts and templates that are specific for parts of the scientific workflow: idea generation, experimentation, and paper writing (lu2024ai). The prompts and templates are currently hand-written and tailored for the specific type of experiment. We are not aware of any meta-science-tools, where these prompts and templates are generated by an LLM.
To help answer the question how well agentic LLM scientists are able to perform, the MLGym framework (nathani2025mlgym) has been introduced. MLGym is designed as an open framework to benchmark AI science tools on a range of domains—from data science, game theory, computer vision, natural language processing, and reinforcement learning—all from machine learning. On a test in 2025 with then-current frontier models, the authors report that prompts and hyperparameter settings are important, and with tuning, results can be improved. They also report that MLGym did not find that models could generate novel hypotheses, algorithms, architectures, or make substantial scientific improvements (nathani2025mlgym). In particular, the authors note that modern LLM agents can successfully tackle a diverse array of quantitative experiments, reflecting advanced skills and domain adaptability but also that it is not yet clear if the notion of scientific novelty can be successfully automated or even formally defined in a form suitable for agents.
Even if current AI science tools are not able to produce breakthrough science results independently, existing agentic LLM tools offer tangible productivity gains for researchers, ranging from literature analysis, idea generation, experiment setup, to improving the writing style. These productivity gains are real, and are having an impact on science. Blogs and videos are changing the way in which ideas are disseminated, and more papers are being written and submitted to conferences and journals, putting pressure on traditional peer review systems. Furthermore, collaborative tools, such as AI Co-scientist, highlight the value of validated research in which agentic tools support human researchers.
3.4.2. Grounding Actions in the Real World
For agents to act in the real world, their understanding must be grounded in the real world. They should sense their surroundings, understand it, and take actions that make sense. LLMs that were only trained on a language corpus may suggest actions such as trying to open doors that do not exist, or moving kitchen items that are not present. World Models and VLAs provide a step towards this world understanding, so that robots and assistants can take actions that make sense.
@@ 3.4.2. Grounding Actions in the Real World
3.4.3. Security, Ethical and Legal aspects of Assistants
@@
Agentic LLMs is an active field of research, some of which is aimed at making assistants ready for commercial deployment. If they work well, there may be a large market for robotic assistants that perform tedious or dangerous work, and for LLM agents that outperform humans in, for example, medical and trading decisions. However, such commercial deployment is still some time into the future, also because important ethical and legal questions should be resolved. If an LLM assistant provides medical advice, and a patient suffers, who is responsible? If an assistant suggests a certain trade, and a trader loses a sum of money, who is liable? Also, the impact on society and the work force has economic implications. Further research is necessary to resolve these questions before assistants can be used in the world in a responsible manner (akata2020research).
4. Interacting
We will now turn to the third category of the survey: interacting agents. Traditional LLMs passively respond to user queries, have no memories of interaction histories beyond their context window, and do not plan future steps of interaction ahead. This is shifting with agentic LLMs: LLMs have memories and planning abilities. Reflective loops can lead to actions at their own initiative. This opens new potential for studying social interaction with users and other machine agents.
In this section we first briefly discuss social and interactive capabilities in traditional, non-agentic LLMs, to identify the roots of their ability to interact with users and agents. Second, we discuss pairs or small teams of agentic LLMs that have role-based interactions to complete a task, game, or experiment. Third, we turn to open-ended interactions of LLM agents, interacting semi-spontaneously without prior role assignment, forming LLM societies that show self-organizing behavior, social dynamics, and emergent norms.
In Section 4.4 we will discuss two influential approaches: CAMEL and Generative Agents, in more detail.
4.1. Social Capabilities of LLMs